# Hamiltonian Active Matter in Incompressible Fluid Membranes

### v1

@@ Abstract
Active proteins and membrane-bound motors exert force dipole flows along fluid interfaces and lipid bilayers. We develop a unified hydrodynamic and Hamiltonian framework for the interactions of pusher and puller dipoles embedded in an incompressible two-dimensional membrane supported by a shallow viscous subphase. Beginning from the screened Stokes equations of the membrane–subphase composite, we derive the real-space incompressible Green’s tensor, obtain its near- and far-field asymptotics, and construct the resulting dipolar velocity and stream functions. Although generic dipoles reorient under the local membrane vorticity, we show that the far-field dipolar flow is vorticity-free; force-free motors therefore retain fixed orientation and obey a Hamiltonian dynamics in which the positions of $N$ dipoles evolve via an effective Hamiltonian built from the dipolar stream function. In the near field, where the flow possesses finite vorticity, a Hamiltonian formulation is recovered in the quenched-orientation limit. Exploiting this structure, we simulate ensembles of pusher and puller dipoles and compare the dynamics generated by the $1/r$ near-field kernel and the subphase screened $1/r^{3}$ far-field kernel. For identical dipoles, the far-field Hamiltonian produces rapid clustering from random initial conditions, whereas the near-field Hamiltonian suppresses collapse and yields extended, non-aggregating configurations.
@@ IV Hamiltonian formulation for dipoles
Figure 3: Representative far–zone dynamics of a randomly initialized cluster of twelve pushers in the incompressible membrane with random initial locations (within a disc) and random orientations. Green disks indicate initial positions, red disks indicate final positions, and the grey curves trace the particle trajectories.
@@ IV.2 Far–field Hamiltonian
generate the $1/r^{3}$ screened flows of Eq. (17). For co–aligned dipoles, both Hamiltonians reduce to simple functions of the relative polar angles, $H\propto\sum_{i\neq j}\sigma_{i}\sigma_{j}\sin(2\theta_{ij})/r_{ij}^{n}$ with $n=0$ or $2$.
@@ V.1 Simulation details
We integrate the $2N$ position variables $\{x_{i}(t),y_{i}(t)\}$ using a standard adaptive Runge–Kutta method, evolving Eq. (25) with either $H_{\rm near}$ or $H_{\rm far}$. Throughout this section we take $N=15$ and use identical random initial positions for both kernels: particles are sampled uniformly from a unit disc. Initial positions are indicated by open circles in Fig. 2, while filled black circles mark the final positions at $t=t_{\rm max}$.
To quantify aggregation or dispersion, we monitor the mean pairwise separation
@@
The similarity between the pusher and puller panels highlights a central result of this work: the nature of the hydrodynamic kernel, rather than the motor type alone, determines whether active dipoles aggregate or disperse. The far–field screened flow is irrotational and generates an attractive effective Hamiltonian, while the near–field unscreened flow frustrates collective collapse.
For completeness, Fig. 3 shows a representative far–zone evolution with arbitrary orientations, illustrating that the same aggregation mechanism persists beyond the co–aligned limit.

### v2

@@ Abstract
Active proteins and membrane-bound motors exert force dipole flows along fluid interfaces and lipid bilayers. We develop a unified hydrodynamic and Hamiltonian framework for the interactions of pusher and puller dipoles embedded in an incompressible two-dimensional membrane supported by a shallow viscous subphase. Beginning from the screened Stokes equations of the membrane–subphase composite, we derive the real-space incompressible Green’s tensor, obtain its near- and far-field asymptotics, and construct the resulting dipolar velocity and stream functions. Although generic dipoles reorient under the local membrane vorticity, we show that the far-field dipolar flow is vorticity-free; force-free motors therefore retain fixed orientation and obey a Hamiltonian dynamics in which the positions of NN dipoles evolve via an effective Hamiltonian built from the dipolar stream function. In the near field, where the flow possesses finite vorticity, a Hamiltonian formulation is recovered in the quenched-orientation limit. Exploiting this structure, we simulate ensembles of pusher and puller dipoles and compare the dynamics generated by the $1/r$ near-field kernel and the subphase screened $1/r^{3}$ far-field kernel. For identical dipoles, the far-field Hamiltonian produces rapid clustering from random initial conditions, whereas the near-field Hamiltonian suppresses collapse and yields extended, non-aggregating configurations.
@@ IV Hamiltonian formulation for dipoles
Figure 3: Representative far–zone dynamics of a randomly initialized cluster of twelve pushers in the incompressible membrane with random initial locations (within a disc) and random orientations. Green disks indicate initial positions, red disks indicate final positions, and the grey curves trace the particle trajectories. A soft harmonic repulsion is included to regularize close encounters.
@@ IV.2 Far–field Hamiltonian
generate the $1/r^{3}$ screened flows of Eq. (17). For co–aligned dipoles, both Hamiltonians reduce to simple functions of the relative polar angles, $H\propto\sum_{i\neq j}\sigma_{i}\sigma_{j}\sin(2\theta_{ij})/r_{ij}^{n}$ with $n=0$ or 22.
@@ V.1 Simulation details
We integrate the $2N$ position variables $\{x_{i}(t),y_{i}(t)\}$ using a standard adaptive Runge–Kutta method, evolving Eq. (25) with either $H_{\rm near}$ or $H_{\rm far}$. Throughout this section we take $N=15$ and use identical random initial positions for both kernels: particles are sampled uniformly from a unit disc. Initial positions are indicated by open circles in Fig. 2, while filled black circles mark the final positions at $t=t_{\rm max}$. To quantify aggregation or dispersion, we monitor the mean pairwise separation
@@
The similarity between the pusher and puller panels highlights a central result of this work: the nature of the hydrodynamic kernel, rather than the motor type alone, determines whether active dipoles aggregate or disperse. The far–field screened flow is irrotational and generates an attractive effective Hamiltonian, while the near–field unscreened flow frustrates collective collapse. For completeness, Fig. 3 shows a representative far–zone evolution with arbitrary orientations, illustrating that the same aggregation mechanism persists beyond the co–aligned limit. Representative far- and near-zone trajectories of co-aligned pusher and puller clusters are provided in the Supplementary Material (Figs. S1–S2). In all these examples, a soft harmonic repulsion is added to regularize close encounters and suppress particle overlap.
@@ Acknowledgements.
Supplemental Material for
“Hamiltonian Active Matter in Incompressible Fluid Membranes”
(a) Pushers.
(b) Pullers.
Figure S1: Far-zone dynamics of co-aligned clusters of twelve pushers (left) and pullers (right) in an incompressible membrane, initialized at random locations within a disc. Initial positions appear in green (pushers) and yellow (pullers), final positions in red and orange, with trajectories in grey. A soft harmonic repulsion is included to prevent particle overlap.
(a) Pushers.
(b) Pullers.
Figure S2: Representative near-zone dynamics of randomly initialized (within a disc), co-aligned clusters of twelve pushers (left) and pullers (right) in an incompressible membrane. Initial positions are shown in green (pushers) and yellow (pullers); final positions in red and orange, respectively; trajectories are grey. A soft harmonic repulsion is included to regularize close encounters.